OpenStudy (anonymous):

What is the equation of the line that passes through the point (3, 2) and has a slope of 4?

OpenStudy (aroub):

y=mx+b , m=4 , (3,2) 2=4(3)+b 2=12+b -10=b =>y=4x-10